OpenStudy (anonymous):

What is the percent increase if 250 guitars used to be made a month but now 15,000 are made a month

OpenStudy (anonymous):

Dive what they made the following month (15,000) by the previous one (250) Then convert from decimal point to percentage by moving the decimal point two times to the right.

OpenStudy (anonymous):

\[\frac{15000}{250}=?\]

OpenStudy (anonymous):

Oh ok I did it a way differnet way

OpenStudy (anonymous):

can you post it up the way you did it?

OpenStudy (lgbasallote):

you can also use the formula \[\large \mathtt{\frac{\text{final money} - \text{initial money}}{\text{initial money}}}\]

OpenStudy (anonymous):

I did 15000-250=14750 then 14750/15000

OpenStudy (anonymous):

I'm sure my answer isn't correct so which format do I use to answer this

OpenStudy (anonymous):

i think the second part you are suppose to divide by 250 not 15000 According to @lgbasallote

OpenStudy (anonymous):

So i got 59%

OpenStudy (anonymous):

i think the answer is 6000% it does not make sense to have the answer be less than 100% an answer of less than 100% implies that there was a decrease

OpenStudy (anonymous):

haha now im confused

OpenStudy (anonymous):

well, 50% = .50, which is less than 1, which implies that there was a decrease. in order for there to be an increase, you want a number larger than 1. which means you want a number over 100% Think about it - what is 50% of 4. Will you get a larger or smaller number?

OpenStudy (anonymous):

50% = 1/2 = .5, remember

OpenStudy (anonymous):

yes i understand that

OpenStudy (anonymous):

How would I put this in a equation?

OpenStudy (anonymous):

so then, in order to get a number larger than 1 (in this case), you have to divide (new amount)/(old amount) then multiply by 100 to get a percent

OpenStudy (anonymous):

but your percent will be 6000%, it seems unnatural cause it is such a large increase, but yeah, 'tis the answer

OpenStudy (anonymous):

wouldn't it be 600?

OpenStudy (anonymous):

oh no no sorry your right

OpenStudy (anonymous):

you can check the answer by going (old amount * percent increase)/100 you should end up with the new amount again (it's the inverse equation)

OpenStudy (anonymous):

thank you ;)